OpenStudy (anonymous):

The volume of a cylinder is V = r2h. Solve for h (height) in this formula.

OpenStudy (anonymous):

do you mean v=(r^2)h?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

To solve for h just get it by itself. (V/r^2) =h

OpenStudy (anonymous):

divide both sides by r^2

OpenStudy (anonymous):

so what the answer ?

OpenStudy (anonymous):

I said it above. (V/r^2) =h

OpenStudy (anonymous):

Unless you have variables for V and r

OpenStudy (anonymous):

its says its wrong ?

OpenStudy (anonymous):

V = πr2h.

OpenStudy (anonymous):

does it give you values for v and r?

OpenStudy (anonymous):

pi!!!!! But you never mentioned pi mariisol.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i just saw it -_-

OpenStudy (anonymous):

lol, well at least I know it's not my fault. xD

OpenStudy (anonymous):

so divide both sides by pi and r^2, giving you h=(v/(pi(r^2)))

OpenStudy (anonymous):

I think mariisol got the problem wrong.